OpenStudy (anonymous):

let f(x)=8x-2, find the equation of the secant line containing (5,f(5) and (8, f(8)

OpenStudy (anonymous):

it's means line containing: (5,38) & (8,62) f(5)=8*5-2=38 f(8)=8*8-2=62

OpenStudy (anonymous):

m=(62-38)/(8-5)=24/3=8 the equation of the line: y-38=8(x-5) y=8x-2